OpenStudy (anonymous):

With a certain tail wind a jet aircraft arrives at its destination, 1,890 miles away, in 3 hours. Flying against the same wind, the plane makes the return trip in 3 3/8 hours. Find the wind speed and the plane's airspeed. plz help!!

OpenStudy (anonymous):

Variables: d = distance r = rate t = time c = speed of current Given: 1,890 miles away Means: d = 1890 Given: trip with tail wind take 3 hours Means: t_tailwind = 3 Given: Flying against the same wind, the plane makes the return trip in 3 3/8 hours Means: t_headwind = 3 3/8 Convert 3 3/8 hours to a decimal. 3 3/8 hours = 3 hours + 3/8 hours = 3 hours + 0.375 hours = 3.375 hours HINT: Write what you know in a table. ..................... distance ....... rate ...... time head wind ...... 1890 ........... r - w ...… 3.375 tail wind …...... 1890 ........... r + w ..… 3 total ............... 3780 ........... ---- ........ 6.375 Remember that the distance formula is d = r * t. Apply this to the table. ...................... distance ................... rate ...... time head wind ...... 1890 = 3.375(r - w) .... r - w ..… 3.375 tail wind …...... 1890 = 3(r + w) ......... r + w .… 3 total ............... 3780 ....................... ---- ........ 6.375 You have 2 equations to use. 1890 = 3.375(r - w) 1890 = 3(r + w) Solve the first one for r. 1890 = 3.375(r - w) 1890 = 3.375(r) + 3.375(-w) 1890 = 3.375r - 3.375w 1890 + 3.375w = 3.375r (1890 + 3.375w) / 3.375 = 3.375r / 3.375 (1890 / 3.375) + (3.375w / 3.375) = r 560 + w = r Plug this into the second equation. 1890 = 3(r + w) 1890 = 3(560 + w + w) 1890 = 3(560 + 2w) 1890 = 3(560) + 3(2w) 1890 = 1680 + 6w 1890 - 1680 = 6w 1890 - 1680 = 6w 210 = 6w 210 / 6 = 6w / 6 35 = w Plug this into 560 + w = r to find r. 560 + w = r 560 + 35 = r 595 = r ANSWER: The wind speed is 35 mph; the plane's airspeed is 595 mph. CHECK: ...................... distance ....................... rate ............ time head wind ...... 1890 = 3.375(595 - 35) .... 595 - 35 ..… 3.375 tail wind …...... 1890 = 3(595 + 35) ......... 595 + 35 .… 3 total ............... 3780 ............................. ---- ............. 6.375 ...................... distance .............. rate ..... time head wind ...... 1890 = 1890 ......... 560 ..… 3.375 tail wind …...... 1890 = 1890 ......... 630 .… 3 total ............... 3780 .................... ---- ..... 6.375

OpenStudy (perl):

good job, but it would be nice if you give the poster a chance to reply , ie, dont give the complete answer

OpenStudy (anonymous):

okay next time i will do that

OpenStudy (anonymous):

@cierra.whisper thank you cierra